# 1 5 Determine the convergence or divergence of s, qdx. Find if conver- gent. -5 #2 3 Find lim n+ n4 + n3
Answer #1

SOLUTION 5 #1. 27 dx we have f(xl 2x x² - 4 76x) = - 2x - f(1) H(-x) = f(x) f(x) i an odd function 5 => flos da 5 2x x²-4 da